Roll Corporation has 20,000 shares of common stock outstanding. It's financed entirely with equity. The un-levered cost of capital is 14%. The company distributes all of its earnings to equity holders as dividends at the end of each year. And Roll Corporation is subject to a corporate tax rate of 30%.

Roll Corporation estimates that its annual EBIT will be as follow: 20% Bust with EBIT 2600, 50% Expected with EBIT 3620, 30% Expansion with EBIT 4900.

The firm expects the earnings will be unchanged in perpetuity.

A. What is the market value of Roll Corporation?

B. Suppose Roll Corporation issues $8000 of perpetual debt and use the proceeds to repurchase common stock. The bond will sell at par with a 6% annual coupon rate. After the sales of the bonds, Roll Corporation will remain the new capital structure indefinitely.

B1: What is the new value of the firm?

B2. What is the new vale of the firm's equity?

B3. What is the required return on Roll Corporation's equity after restructuring?

B4. Will the presence of corporate tax increase or decrease the firm value? Why?
Answer #1

afected EBIT = a + 4900 x30/ = 520 +1810+1470-3800 20,000 equial gharer unlevered finm uitち Tar 외ala(0-30 / Marikit value o Cost o Capital 3800 x (1 0.3 014 19000 debt 8000 : Cost% dart-Coupon alex(1-t) 6.1.x6-0. 3) τ : Cost o●a00, 000 Adt = $8000 二New eguit_ 3200,000-8000.= 192000 4위, x 8000 = a08000 = 0.00162+0.1344 (B) Valu tfhm = (EBIT-1)s(-t)Assumption: Value per share is $10.

If share value is different answer will modify accordingly